OpenStudy (anonymous):

A 210 second song is divided into 3 parts, the beginning the middle and the end. The beginning is one-sixth as long as the end. The middle is one half as long as the end. Find the length for the beginning the middle and the end.

OpenStudy (dumbcow):

1/6x +1/2x + x = 210 1/6x + 3/6x + 6/6x = 210 10/6x = 210 x = 210*6/10 x = 126